OpenStudy (anonymous):

Original Mass: 28.39g Final Mass: 27.11g What is the percent of change? Please show work

OpenStudy (anonymous):

[ (28.39-27.11) / 28.39 ] * 100 So we get [ 1.28 / 28.39 ] * 100 and this is ~0.0451 * 100 So the percent of change is 4.5%

OpenStudy (anonymous):

roughly

OpenStudy (anonymous):

28.39-27.11=1.28.....1.28/28.39*100=4.5%